# HSR-Enhanced Sparse Attention Acceleration

### v1

@@ 4 Main Results on Attention Generation
3:     HalfSpaceReport hsr $\triangleright$ Algorithm 1, Part 2 of Corollary 3.5
@@
11:     hsr.Init($\{K_{i}\}_{i\in[n]},n,d$) $\triangleright$ It takes ${\cal T}_{\mathsf{init}}(n,d)$ time
@@
16:         $\widetilde{S}_{i,\mathrm{fire}}\leftarrow\textsc{hsr}.\textsc{Query}(Q_{i},b)$triangleright It takes ${\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})$ time
@@
We begin with introducing our result on ReLU attention generation as follows:
@@
Theorem 4.1 (Running time of ReLU attention generation, informal version of Theorem C.2).
@@
Let ReLU attention be defined as Definition 1.2. Assume each entry of $K$ is from Gaussian ${\cal N}(0,\sigma_{k}^{2})$, and each entry of $Q$ is from Gaussian ${\cal N}(0,\sigma_{q}^{2})$. Let $\delta\in(0,1)$ denote the failure probability. Let $\sigma_{a}=4\cdot(1+d^{-1}\log(m/\delta))^{1/2}\cdot\sigma_{q}\sigma_{k}$. Let $b=\sigma_{a}\cdot\sqrt{0.4\log n}$. Suppose we have KV Cache $K,V\in\mathbb{R}^{n\times d}$. We want to generate a $m$ length answer, where $n\gg m$. Then, our inference function in Algorithm 2, with probability at least $1-\delta$, takes $O(mn^{4/5})$ time to generate the answer.
@@
Theorem 4.1 shows that our Algorithm 2 accelerates the running time of ReLU attention generation from naive $O(mn)$ to $O(mn^{4/5})$, which is a significant speed up when the KV Cache is large. The at least $1-\delta$ success probability originates from the sparsity analysis of ReLU attention (Lemma 6.1), where with probability at least $1-\delta$, we have the number of non-zero entries of each row of the attention matrix is at most $n^{4/5}$.
@@
Then, we move on to presenting our result on Softmax attention generation. Our results consist two parts: the improved running time of Softmax attention generation, and the error analysis of Softmax attention with index set. Firstly, we introduce our result about the imporved running time of Softmax attention generation as follows:
@@
Theorem 4.2 (Running time of Softmax attention generation, informal version of Theorem E.1).
@@
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 2 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we can show that the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under the Softmax attention generation scenario: Suppose we have KV Cache $K,V\in\mathbb{R}^{n\times d}$. We want to generate a $m$ length answer, where $n\gg m$. Our inference function in Algorithm 2 (replacing ReLU attention with Softmax attention) takes $O(mn^{4/5})$ time to generate the answer.
@@
Theorem 4.2 demonstrates that if we choose the threshold $b$ satisfying $R=\mathsf{NN}(n^{4/5},q,K)$, we can achieve a significant running time improve of the Softmax attention generation.
@@
It is evident that this method introduces an approximation error due to the exclusion of certain entries. Nevertheless, under mild assumptions about the distribution of the attention scores, we demonstrate that this approximation error is indeed negligible. The proof’s intuitive explanation lies in the fact that the majority of attention scores are focused on the small subset of entries that we retain. We organize our result as follows:
@@
Theorem 4.3 (Error analysis of Softmax attention with index set, informal version of Theorem F.2).
@@
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $q\in\mathbb{R}^{d}$ denote a single row of $Q\in\mathbb{R}^{m\times d}$. Let $\gamma\in[0,1]$, $\beta_{1}\geq\beta_{2}\geq 0$. Let the index set $R$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ denote the indices of top-$r$ entries of $qK$. Let $R=\mathsf{NN}(n^{\gamma},q,K)\subseteq[n]$, where $|R|=n^{\gamma}$. Assume the query $q$ and key cache $K$ have $(\gamma,\beta_{1},\beta_{2})$ massive activation property (Definition 3.3). Then, we have
@@
5 Extension on Full Attention Computation
In this section, we extend our results to full attention computation scenario, where the number of queries and keys is proportional, i.e., $m=\Theta(n)$. Essentially, the full attention computation is beneficial in practical applications, particularly within the context of cross-attention computations. For ReLU attention, we leverage Part 1 result of Corollary 3.5 to accelerate the identification of non-zero entries (activated entries). We introduce our result on ReLU attention as follows:
Algorithm 3 Full attention computation
1:data structure FullAttentionComputation $\triangleright$ Lemma 6.3
@@ 5 Extension on Full Attention Computation
3:     HalfSpaceReport hsr $\triangleright$ Algorithm 1, Part 1 of Corollary 3.5
@@
8:     hsr.Init($\{K_{i}\}_{i\in[n]},n,d$) $\triangleright$ It takes ${\cal T}_{\mathsf{init}}(n,d)$ time
@@
11:         $\widetilde{S}_{i,\mathrm{fire}}\leftarrow\textsc{hsr}.\textsc{Query}(Q_{i},b)$triangleright It takes ${\cal T}_{\mathsf{query}}(n,d,\widetilde{k}_{i})$ time.
@@
Theorem 5.1 (Running time of full ReLU attention computation, informal version of Theorem B.2).
@@
In Theorem 5.1, we improve the running time of full ReLU attention computation from $O(n^{2})$ to $O(n^{2-1/\lfloor d/2\rfloor}+n^{1+4/5})$, which is a notable uplift of the running time when $n$ is extremely large.
Then, we present our result on Softmax attention. Intuitively, we use the Part 1 result of Corollary 3.5 to identify those “massive activated” entries (top-$r$ indices) within the attention matrix of Softmax attention, and calculate the Softmax attention with top-$r$ indices. We organize our result as follows:
Theorem 5.2 (Running time of Softmax full attention computation, informal version of Theorem E.2).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 3 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we have the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under full Softmax attention computation scenario: Suppose we have $m=\Theta(n)$. Algorithm 3 (replacing ReLU attention with Softmax attention) takes $O(n^{2-1/\lfloor d/2\rfloor}+n^{1+4/5})$ time to compute the full ReLU attention of $Q,K,V$.
Theorem 5.2 demonstrates our $O(n^{2-1/\lfloor d/2\rfloor}+n^{1+4/5})$ running time on Softmax full attention computation, which improves from naive running time $O(n^{2})$.
@@ 6 Technical Overview
In Section 6.1, we introduce our analysis about the sparsity in the ReLU attention mechanism. In Section 6.2, we present our results of two general attention frameworks. In Section 6.3, we provide our error analysis of Softmax attention with index set. We have shown that with mild assumption on the distribution of attention scores, the error of Softmax attention with index set is negligible.
@@ 6.1 Sparsity Analysis of ReLU Attention
Lemma 6.1 (Sparsity analysis, informal version of Lemma D.3).
@@ 6.2 General Attention Frameworks
First, we introduce our general framework for attention generation computation. Here, we use the Part 1 result of the HSR data structure. As for this framework is designed for the attention generation task, the key matrix $K$ is fixed in each inference. Therefore, in the Init procedure, we initialize the HSR data structure with the key matrix $K$. Then, in each inference, we use the same HSR data structure to answer the query from each row of the query matrix $Q$. We provide the result of this general attention generation framework as follows.
Lemma 6.2 (General attention generation framework, informal version of Lemma C.1).
@@
The general framework for full attention computation is quite different from the previous one. Namely, we choose the Part 2 result of the HSR data structure. Since in each inference, both the query matrix $Q$ and the key matrix $K$ differ from any other inference, we first initialize the HSR data structure with the key matrix $K$. Then for each row of the query matrix $Q$, we query the HSR data structure to find the activated entries.
Lemma 6.3 (General full attention computation framework, informal version of Lemma B.1).
@@ 6.3 Error Analysis of Softmax Attention with Top-$r$ Indices
Calculating the Softmax attention on the “massive actavted” index set will introduce approximation error. In the following Lemma, we analyze the quantity of this approximation error. Here, we use $\alpha$ to denote the summation of all entries activated by $\exp(x)$ function, and we use $\overline{\alpha}$ to denote the summation of those entries which are excluded from “massive activated” index set. We provide the general error bound of Softmax attention with index set as follows.
Lemma 6.4 (General error analysis of Softmax attention with index set, informal version of Lemma F.1).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $q\in\mathbb{R}^{d}$ denote a single row of $Q\in\mathbb{R}^{m\times d}$. Let $\alpha,\overline{\alpha}$ and $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2. Then we have $\|\mathsf{Attn}_{s}(q,K,V)-\widehat{\mathsf{Attn}}_{s}(q,K,V)\|_{\infty}\leq\frac{2\overline{\alpha}}{\alpha}\cdot\|V\|_{\infty}.$
Note that Lemma 6.4 only provides a general error analysis of Softmax attention with index set. Under mild assumption on the distribution of attention scores, we show that this error is actually very small. For more details, please refer to Theorem 4.3.
@@ 7 Experiments
Datasets. To estimate the approximation error of the Softmax attention with “massive activation” entries, we conduct experiments on the PaulGrahamEssays datasets from LLMTest-NeedleInAHaystack [33]. Specifically, for each article in the dataset, we first input $327682^{15}=$ 32768 tokens to the LLMs, then generate $1024$ tokens.
Metric. We evaluate the generation quality by the classical perplexity. Perplexity is defined as the exponentiated average negative log-likelihood of a sequence. If we have a tokenized sequence $X=(x_{0},x_{1},\cdots,x_{N})$, then the perplexity of $X$ is: $\mathrm{Perplexity}(X)=\exp({-\frac{1}{N}\sum_{i=1}^{N}\log p_{\theta}(x_{i}|x_{<i})}),$ where $\log p_{\theta}(x_{i}|x_{<i})$ is the log-likelihood of the $i$-th token conditioned on the preceding tokens. Intuitively, it can be thought of as an evaluation of the model’s ability to predict uniformly among the set of specified tokens in a corpus. Importantly, the tokenization procedure has a direct impact on a model’s perplexity which should be taken into consideration when comparing different models.
Models. To demonstrate the generalization of our approximation error bound, we conducted experiments on three mainstream large models: LLaMA 3.1 8B Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct [51], Mistral Nemo 12B Instruct [^3] https://huggingface.co/mistralai/Mistral-Nemo-Base-2407 [52], and Phi 3.5 Mini 3.8B Instruct [^4] https://huggingface.co/microsoft/Phi-3.5-mini-instruct [2].
Results. The experiments are conducted on the setting discussed in previous paragraphs. We evaluated the performance of three mainstream LLMs using Softmax attention with top-$r$ indices. In particular, we chose $r$ from the set $\{2^{2},2^{4},2^{6},2^{8},2^{10},2^{12},2^{15}\}$. As depicted in Figure 2, a significant increase in the perplexity (drop in performance) of LLMs is observed only when $r$ falls below $2^{4}$. This suggests that the “massive activated” tokens are predominantly found within the top-$2^{4}$ entries. In comparison to the total of $2^{15}$ entries, the “massive activated” entries constitute a relatively minor fraction. The observed results align with our theoretical analysis, confirming that the approximation error of the Softmax attention mechanism with top-$r$ indices is insignificant for larger values of $r$.

### v2

@@ 4 Main Results on Generation Decoding
3:     HalfSpaceReport hsr $\triangleright$ Part 2 of Corollary 3.1
@@
11:     hsr.Init($\{K_{i}\}_{i\in[n]},n,d$) $\triangleright$ It takes $O(n^{\lfloor d/2\rfloor})$ time
@@
16:         $\widetilde{S}_{i,\mathrm{fire}}\leftarrow\textsc{hsr}.\textsc{Query}(Q_{i},b)$triangleright It takes $O(n^{4/5})$ time
@@
We begin with introducing our result on ReLU attention generation decoding as follows:
@@
Theorem 4.1 (Running time of ReLU attention generation decoding, informal version of Theorem D.2).
@@
Let ReLU attention be defined as Definition 1.2. Assume each entry of $K$ conforms Gaussian ${\cal N}(0,\sigma_{k}^{2})$, and each entry of $Q$ conforms Gaussian ${\cal N}(0,\sigma_{q}^{2})$. Let $\delta\in(0,1)$ denote the failure probability. Let $\sigma_{a}=4\cdot(1+d^{-1}\log(m/\delta))^{1/2}\cdot\sigma_{q}\sigma_{k}$. Let $b=\sigma_{a}\cdot\sqrt{0.4\log n}$. Suppose we have KV Cache $K,V\in\mathbb{R}^{n\times d}$. We want to generate a $m$ length answer, where $n\gg m$. Then, our inference function in Algorithm 1, with probability at least $1-\delta$, takes $O(mn^{4/5})$ time to generate the answer.
@@
Theorem 4.1 shows that our Algorithm 1 accelerates the running time of ReLU attention generation decoding from naive $O(mn)$ to $O(mn^{4/5})$, which is a significant speed up when the KV Cache is large. We provide an example for the sparsity with different $n$ in Table 1 in the Appendix. The at least $1-\delta$ success probability originates from the sparsity analysis of ReLU attention (Lemma 6.1), where with probability at least $1-\delta$, we have the number of non-zero entries of each row of the attention matrix no larger than $n^{4/5}$.
@@
Then, we move on to presenting our result on Softmax attention generation decoding. Our results consist of two parts: the improved running time and the approximation error analysis. Firstly, we introduce our result about the improved running time of Softmax attention generation decoding as follows:
@@
Theorem 4.2 (Running time of Softmax attention generation decoding, informal version of Theorem F.1).
@@
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 1 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we can show that the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under the Softmax attention generation scenario: Suppose we have KV Cache $K,V\in\mathbb{R}^{n\times d}$. We want to generate a $m$ length answer, where $n\gg m$. Our inference function in Algorithm 1 (replacing ReLU attention with Softmax attention) takes $O(mn^{4/5})$ time to generate the answer.
@@
Theorem 4.2 demonstrates that if we choose the threshold $b$ satisfying $R=\mathsf{NN}(n^{4/5},q,K)$, we can achieve a significant running time improvement of the Softmax attention generation.
@@
Indeed, this method introduces an approximation error due to the exclusion of certain entries. Nevertheless, under mild assumptions about the distribution of the attention scores, we demonstrate that this approximation error is indeed negligible. The proof’s intuitive explanation lies in the fact that the majority of attention scores are focused on the small subset of entries that we retain. We organize our results as follows:
@@
Theorem 4.3 (Error analysis of Softmax attention with index set, informal version of Theorem G.2).
@@
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $q\in\mathbb{R}^{d}$ denote a single row of $Q\in\mathbb{R}^{m\times d}$. Let $\gamma\in[0,1]$, $\beta_{1}\geq\beta_{2}\geq 0$. Let the index set $R$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ denote the indices of top-$r$ entries of $qK$. Let $R=\mathsf{NN}(n^{\gamma},q,K)\subseteq[n]$, where $|R|=n^{\gamma}$. Assume the query $q$ and key cache $K$ have $(\gamma,\beta_{1},\beta_{2})$ massive activation property (Definition B.3). Then, we have
@@
5 Extension on Prompt Prefilling
In this section, we extend our results to prompt prefilling scenarios where the number of queries and keys is proportional, i.e., $m=\Theta(n)$. For ReLU attention, we leverage Part 1 result of Corollary 3.1 to accelerate the identification of non-zero entries (activated entries). We introduce our result on ReLU attention as follows:
Algorithm 2 Prompt Prefilling
1:data structure PromptPrefilling $\triangleright$ Lemma 6.3
@@ 5 Extension on Prompt Prefilling
3:     HalfSpaceReport hsr $\triangleright$ Part 1 of Corollary 3.1
@@
8:     hsr.Init($\{K_{i}\}_{i\in[n]},n,d$) $\triangleright$ It takes $O(n\log n)$ time
@@
11:         $\widetilde{S}_{i,\mathrm{fire}}\leftarrow\textsc{hsr}.\textsc{Query}(Q_{i},b)$triangleright It takes $O(n^{1-1/\lfloor d/2\rfloor}+\widetilde{k}_{i})$ time.
@@
Theorem 5.1 (Running time of ReLU attention prompt prefilling, informal version of Theorem C.2).
@@
In Theorem 5.1, we improve the running time of ReLU attention prompt prefilling from $O(n^{2})$ to $O(n^{2-1/\lfloor d/2\rfloor}+n^{1+4/5})$, which is a notable uplift of the running time when $n$ is extremely large.
Then, we present our result on Softmax attention. Intuitively, we use the Part 1 result of Corollary 3.1 to identify those “massive activated” entries (top-$r$ indices) within the attention matrix of Softmax attention and calculate the Softmax attention with top-$r$ indices. We organize our results as follows:
Theorem 5.2 (Running time of Softmax attention prompt prefilling, informal version of Theorem F.2).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $\mathsf{NN}(r,q,K)\subseteq[n]$ and the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2. We choose the threshold $b\in\mathbb{R}$ in Algorithm 2 such that $R=\mathsf{NN}(n^{4/5},q,K)$. Then, we have the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ achieves outstanding running time under Softmax attention prompt prefilling scenario: Suppose we have $m=\Theta(n)$. Algorithm 2 (replacing ReLU attention with Softmax attention) takes $O(n^{2-1/\lfloor d/2\rfloor}+n^{1+4/5})$ time to compute the full ReLU attention of $Q,K,V$.
Theorem 5.2 demonstrates our $O(n^{2-1/\lfloor d/2\rfloor}+n^{1+4/5})$ running time on Softmax attention prompt prefilling, which improves from naive running time $O(n^{2})$.
@@ 6 Technical Overview
Section 6.1 introduces our analysis of the sparsity in the ReLU attention mechanism. Section 6.2 presents our results of two general attention frameworks. Section 6.3 provides our error analysis of Softmax attention with index set. We have shown that with mild assumption on the distribution of attention scores, the error of Softmax attention with index set is negligible.
@@ 6.1 Sparsity Analysis of ReLU Attention
Lemma 6.1 (Sparsity analysis, informal version of Lemma E.3).
@@ 6.2 General Attention Frameworks
First, we introduce our general framework for generation decoding. Here, we use the Part 2 result of the HSR data structure. As this framework is designed for the generation decoding task, the key matrix $K$ is fixed in each inference. Therefore, in the Init procedure, we initialize the HSR data structure with the key matrix $K$. Then, in each inference, we use the same HSR data structure to answer the query from each row of the query matrix $Q$. We provide the result of this general generation decoding framework as follows.
Lemma 6.2 (General generation decoding framework, informal version of Lemma D.1).
@@
The general framework for prompt prefilling is quite different from the previous one. Namely, we choose the Part 1 result of the HSR data structure. Since in each inference, both the query matrix $Q$ and the key matrix $K$ differ from any other inference, we first initialize the HSR data structure with the key matrix $K$. Then, for each row of the query matrix $Q$, we query the HSR data structure to find the activated entries.
Lemma 6.3 (General prompt prefilling framework, informal version of Lemma C.1).
@@
Then, we use the following Remark to demonstrate the different intuitions behind Lemma 6.2 (Algorithm 1) and Lemma 6.3 (Algorithm 2).
Remark 6.4.
Algorithm 1 is tailored for generation decoding scenarios, where the key matrix $K$ remains constant throughout each inference. Consequently, our optimization efforts are directed at decreasing the time required for individual inferences, which is achieved by adopting Part 2 of Corollary 3.1. In contrast, Algorithm 2 is intended for prompt prefilling, a context in which the key matrix $K$ varies with each inference. Thus, our objective shifts to minimizing the time complexity associated with initializing the HSR data structure, leading us to select Part 1 of Corollary 3.1. For more details, please refer to Figure 2.
@@ 6.3 Error Analysis of Softmax Attention with Top-$r$ Indices
Calculating the Softmax attention on the “massive activated” index set will introduce an approximation error. In the following Lemma, we analyze the quantity of this approximation error. Here, we use $\alpha$ to denote the summation of all entries activated by $\exp(x)$ function, and we use $\overline{\alpha}$ to denote the summation of those entries which are excluded from “massive activated” index set. We provide the general error bound of Softmax attention with an index set as follows:
Lemma 6.5 (General error analysis of Softmax attention with index set, informal version of Lemma G.1).
Let $Q\in\mathbb{R}^{m\times d}$, $K,V\in\mathbb{R}^{n\times d}$ and the Softmax attention $\mathsf{Attn}_{s}$ be defined in Definition 1.1. Let $q\in\mathbb{R}^{d}$ denote a single row of $Q\in\mathbb{R}^{m\times d}$. Let $\alpha,\overline{\alpha}$ and $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2. Then, we have
$\|\mathsf{Attn}_{s}(q,K,V)-\widehat{\mathsf{Attn}}_{s}(q,K,V)\|_{\infty}\leq\frac{2\overline{\alpha}}{\alpha}\cdot\|V\|_{\infty}.$
Note that Lemma 6.5 only provides a general error analysis of Softmax attention with an index set. Under mild assumption on the distribution of attention scores, we show that this error is actually very small. For more details, please refer to Theorem 4.3.
@@ 7 Experiments
Datasets.
To estimate the approximation error of the Softmax attention with “massive activation” entries, we conduct experiments on the PaulGrahamEssays datasets from LLMTest-NeedleInAHaystack [42]. Specifically, for each article in the dataset, we first input $327682^{15}=$ 32768 tokens to the LLMs, then generate $1024$ tokens.
Metric.
We evaluate the generation quality by the classical perplexity. Perplexity is defined as the exponentiated average negative log-likelihood of a sequence. If we have a tokenized sequence $X=(x_{0},x_{1},\cdots,x_{N})$, then the perplexity of $X$ is: $\mathrm{Perplexity}(X)=\exp({-\frac{1}{N}\sum_{i=1}^{N}\log p_{\theta}(x_{i}|x_{<i})}),$ where $\log p_{\theta}(x_{i}|x_{<i})$ is the log-likelihood of the $i$-th token conditioned on the preceding tokens. Intuitively, it can be thought of as an evaluation of the model’s ability to predict uniformly among the set of specified tokens in a corpus. Importantly, the tokenization procedure has a direct impact on a model’s perplexity, which should be taken into consideration when comparing different models.
Models.
To demonstrate the generalization of our approximation error bound, we conducted experiments on three mainstream large models: LLaMA 3.1 8B Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct [65], Mistral Nemo 12B Instruct [^3] https://huggingface.co/mistralai/Mistral-Nemo-Base-2407 [66], and Phi 3.5 Mini 3.8B Instruct [^4] https://huggingface.co/microsoft/Phi-3.5-mini-instruct [2].
Results.
The experiments are conducted in the setting discussed in previous paragraphs. We evaluated the performance of three mainstream LLMs using Softmax attention with top-$r$ indices. In particular, we chose $r$ from the set $\{2^{2},2^{4},2^{6},2^{8},2^{10},2^{12},2^{15}\}$. As depicted in Figure 3, a significant increase in the perplexity (drop in performance) of LLMs is observed only when $r$ falls below $2^{4}$. This suggests that the “massive activated” tokens are predominantly found within the top-$2^{4}$ entries. In comparison to the total of $2^{15}$ entries, the “massive activated” entries constitute a relatively minor fraction. The observed results align with our theoretical analysis, confirming the approximation error of the Softmax attention with top-$r$ indices is negligible for larger values of $r$.